OpenStudy (anonymous):

3/5 divided by 12/25 fractions? & could you explain how to get it ! please&thanks in advance.

OpenStudy (dumbcow):

Flip and multiply \[\frac{3}{5} *\frac{25}{12} = \frac{5}{4}\]

OpenStudy (anonymous):

you only flipped twenty five twelfts . & i dont understand how you got five fourths...did you simplify?

OpenStudy (dumbcow):

sorry, yes only flip the bottom fraction i just reduced 3/12 = 1/4 25/5 = 5

OpenStudy (anonymous):

thanks, so the second one is the reciprocol right ?